# IF-Critic: Towards a Fine-Grained LLM Critic for Instruction-Following Evaluation

### v1

@@ Introduction
After obtaining the SFT critic, we seek to further enhance its capability via preference optimization. Existing methods usually sample multiple independent responses from the policy model and utilize response-level rewards to select chosen and rejected examples (Yuan et al. 2024; Ye et al. 2025b; Yu et al. 2025). However, since the critiques in our task involve multiple segments for different constraints in the checklist, only the segments where chosen and rejected critiques differ in judgments encapsulate crucial preference information. Other segments may introduce unnecessary interference and dilute important supervision signals, leading to insufficient optimization (Chen et al. 2024; Cheng et al. 2025).
To mitigate this, we introduce fine-grained, constraint-level comparisons in preference pairs construction, thus reinforcing the perception of crucial preference information. Specifically, for each instruction $x$ in $D_{\text{ref}}$, we first sample $M$ critiques from the SFT critic. Then, we identify the best self-sampled explanation $\hat{e}_{k}$ for each constraint $c_{k}$ whose judgements align with the expert final judgement $j^{*}_{k}$ in a manner analogous to Equation 1 and 2 (if any). Finally, self-sampled critiques with at least one judgment that misaligns with the expert final critique are selected as rejected critiques $C_{l}=\bigcup_{k=1}^{n}(e_{k},j_{k})$. Retaining the segments that align with the expert final critique while replacing other misaligned segments with the best self-sampled explanations $\hat{e}_{k}$ to construct chosen critique $C_{w}$:
@@
Rejected critiques are discarded when some misaligned segments lack substitute best self-sampled explanations. Consequently, we construct a preference dataset $D_{\text{dpo}}$ where the preference pairs only differ in the segments with inconsistent judgments, thereby reinforcing crucial preference information. Moreover, self-sampled critiques are closer to the SFT critic decoding space than the expert critique, which is generally beneficial for optimization (Guo et al. 2024). Finally, we conduct preference optimization via the following DPO loss:
@@
Evaluation Results Utilization
@@
Experiments
Instruction-Following Evaluation
@@
Considering the scarcity of existing instruction-following meta-evaluation datasets, we sample instructions from four benchmarks: EvalCritic, CFBench (Zhang et al. 2024a), TRACE (Zhang et al. 2024c), and Multi-IF (He et al. 2024b) to construct meta-evaluation datasets. EvalCritic is a random split of the dataset we used for instruction-following optimization [^2] We elaborate on this dataset in the following ”Instruction-Following Optimization” section., whose distribution of instructions is distinct from the training data of IF-Critic. For the first three benchmarks, we randomly sample 150 instructions. For Multi-IF, we select the first round from the Chinese version, comprising 454 instructions. Responses for each instruction are generated by 2 out of 6 LLMs with varying capabilities. We use our checklist generator to generate constraint checklists for EvalCritic and directly adapt the original checklists for other benchmarks. For the first three benchmarks, two annotators are required to independently judge the following of each constraint, with a third inspector conducting cross-validation. Any discrepancies will be discussed to reach a consensus. For Multi-IF, we use the provided verification code for evaluation. Across these benchmarks, there are 1411 / 1070 / 1634 / 1220 constraints, respectively, of which 406 / 252 / 297 / 320 are labeled as ”not followed”.
Prompt Model EvalCritic CFBench TRACE Multi-IF (zh) Avg.
@@
Constraint-Level o4-mini 0.887 0.717 0.927 0.755 0.923 0.599 0.960 0.890 0.832
@@
Checklist-Level o4-mini 0.911 0.753 0.935 0.761 0.925 0.639 0.964 0.900 0.849
@@
We choose state-of-the-art general and evaluation-specific LLMs as our baselines. (1) General LLMs: We adopt 4 advanced proprietary LLMs, GLM-4-Plus (GLM et al. 2024), GPT-4.1 (OpenAI 2023), Deepseek-R1 (Guo et al. 2025a) and o4-mini (Jaech et al. 2024), as well as 4 strong open-source LLMs, GLM-4-32B-0414 (GLM et al. 2024), QwQ-32B (Team 2025), LLama-3.3-70B-Instruct (Grattafiori et al. 2024), and Qwen2.5-72B-Instruct (Yang et al. 2024) as our general baselines. These models can serve as evaluators for instruction-following via elaborate prompts without further training. We explore two prompting strategies: 1) Checklist-Level Prompt: requires the LLMs to evaluate the following of all constraints in the given checklist within a single inference pass, which is the same as IF-Critic. 2) Constraint-Level Prompt: requires the LLMs to evaluate each constraint individually. (2) Evaluation-specific LLMs: We select Skywork-Reward-V2-Llama-3.1-8B-40M (Liu et al. 2025a), Prometheus-BGB-8x7B (Kim et al. 2024b), RM-R1-Deepseek-Distilled-Qwen-32B (Chen et al. 2025), and RRM-32B (Guo et al. 2025b) as our evaluation-specific baselines. These models are mainly designed for pairwise comparison between two responses. To accommodate this evaluation setting, we convert the above datasets into pairwise versions. Specifically, for two responses generated from the same instruction, we select pairs in which one response follows all the constraints while the other does not to construct pairwise samples. The rewards determined by the critiques of IF-Critic according to Equation 6 for both responses are converted into pairwise comparisons to calculate the agreement rate. To ensure a fair comparison, we remove any samples that result in a tie for IF-Critic. Greedy search decoding is used for reproducibility.
@@
We choose Qwen2.5-14B-Instruct (Yang et al. 2024) as the base model for both the checklist generator and IF-Critic. Deepseek-R1 is required to generate $N=5$ expert critiques per response. We use difflab [^3] https://docs.python.org/3/library/difflib.html to implement the text similarity measure $u(\cdot)$ for final explanation selection. During IF-Critic training, the ratio of $D_{\text{sft}}$ to $D_{\text{ref}}$ is set to 6:4. The number of self-sampled critiques, $M$, is set to 10, and up to 1 preference pair is constructed for each evaluation input. More details of IF-Critic training are in the Appendix.
@@
Table 3: The average of Positive F1 and Negative F1 under different ablation settings for IF-Critic.
@@
Table 4: Experimental results (%) on Multi-IF, CFBench, and SysBench. Relative Time refers to the relative time to IF-Critic of reward calculation in DPO and average per-step training in GRPO. Skywork-Reward-V2-Llama-3.1-8B-40M is a reward model that generates scalar rewards, whereas QwQ-32B and IF-Critic generate natural language critiques.
@@
Table 1 presents the F1-scores for both positive and negative classes in the instruction-following evaluation task of general LLMs, while Table 2 shows the pairwise agreement for evaluation-specific LLMs. Firstly, IF-Critic demonstrates superior performance on various evaluation benchmarks, outperforming all baselines on average F1. Notably, it surpasses the expert critiques generation models, Deepseek-R1, on all benchmarks, and o4-mini on three out of four benchmarks. In contrast, the performance of existing general LLMs remains limited, with relatively low negative F1. This aligns with our motivation for developing IF-Critic to enable more reliable evaluation. Secondly, existing evaluation-specific LLMs still perform poorly in evaluating instruction-following and lag far behind IF-Critic, highlighting the necessity of a critic specifically designed for this task. Finally, reasoning models (i.e., o4-mini, Deepseek-R1, and QwQ-32B) perform better with Checklist-Level Prompt, while other LLMs benefit more from Constraint-Level Prompt. We attribute this to two factors: (1) Evaluating multiple constraints at one inference pass increases task complexity. (2) With the assistance of the constraint checklist, long-chain reasoning enables a more holistic perception of the relationships between constraints. This validates the rationale of our checklist-guided critique generation paradigm.
@@
We conduct ablation studies on three main factors of our framework: training data format, critique filtering mechanism, and preference optimization method. For the training data format, we decompose the expert critiques to train an alternative model that generates critiques for individual constraints. Table 3 shows that the performance of IF-Critic declines across all benchmarks, demonstrating that checklist-guided training can facilitate the perception of instructions and contribute to the final performance.
For the critique filtering mechanism, we remove each step individually to explore its impact on evaluation performance. Table 3 shows that all steps contribute to the final performance, with the removal of rule-augmented verification yielding the largest performance drop, underscoring its key role in improving the discriminability for length constraints. Notably, removing the entire mechanism and directly using all raw training data leads to the most severe performance decline, highlighting the effectiveness of critique filtering.
For the preference optimization method, we explored three variants of our method: (1) Constructing preference pairs directly based on the agreement rates between the expert critiques, denoted as Vanilla DPO; (2) Replacing misaligned segments of rejected critiques with expert critiques rather than the best self-sampled critiques; and (3) Performing SFT on the entire critique training dataset without preference optimization. Table 3 demonstrates that our constraint-level preference optimization method achieves the best performance, evidencing its effectiveness in enabling better alignment with expert critiques.
@@
To assess the quality of the explanations generated by IF-Critic, we randomly sample 300 constraints from the first three evaluation benchmarks and collect corresponding generated explanations from IF-Critic, QwQ-32B, Deepseek-R1, and o4-mini. For each pair of explanations (one from IF-Critic and the other from a baseline evaluation model, given the same evaluation input), we ask human annotators to judge which explanation is better (i.e., win, lose, or tie) in terms of correctness, informativeness, and helpfulness. The priority of these aspects follows the above order. As shown in Figure 3, IF-Critic surpasses QwQ-32B and Deepseek-R1 in win rates by 9.3% and 7.7%, respectively, and performs on par with o4-mini. This demonstrates that IF-Critic can generate high-quality explanations for instruction-following evaluation.
Instruction-Following Optimization
@@
To collect diverse and high-quality complex instructions for model training, we first curate a set of seed instructions from real application scenarios. These scenarios are distinct from the training data of IF-Critic and generally exhibit lower instruction complexity. Stratified sampling based on task categories and LLM-based quality filtering is adopted to enhance their diversity and quality. Subsequently, we employ Deepseek-R1 to introduce constraints to these seed instructions, guided by a comprehensive constraint taxonomy encompassing 6 primary constraint categories and 4 constraint composition types, thereby increasing their complexity. Deepseek-R1 is also leveraged to validate the generated instructions, ensuring they are reasonable, unambiguous, and logically consistent. Further details are in the Appendix. In total, we obtain 18.5k high-quality complex instructions for training.
We choose Qwen2.5-7B-Instruct and Llama-3.1-8B-Instruct as the initial policy models for DPO and GRPO training. Three evaluation models are used to provide reward signals: IF-Critic, the state-of-the-art reward model Skywork-Reward-V2-Llama-3.1-8B-40M, and open-source general LLM QwQ-32B. For DPO training, we sample 10 responses per instruction, with a temperature of 1.0 and a top-p value of 0.9. The responses with the highest and lowest rewards are used to construct preference pairs. For GRPO training, we perform 32 rollouts per instruction for value estimation. More training details can be found in the Appendix.
@@
We evaluate our models on three popular instruction-following benchmarks: Multi-IF (He et al. 2024b), which assesses multi-turn and multilingual hard constraints following; CFBench (Zhang et al. 2024a), which covers a comprehensive range of constraint types; and Sysbench (Qin et al. 2025), which focuses on the following of system prompts. GPT-4o-2024-11-20 is used as the evaluation model for CFBench and SysBench. For Multi-IF, we follow RPO (Huang et al. 2025) and report the average performance across the English and Chinese subsets.
@@
The main results are presented in Table 4. Firstly, optimization with IF-Critic brings substantially superior performance gains across various benchmarks compared to all strong baselines, substantiating its advanced capacity of reliable instruction-following evaluation. Notably, LLMs after GRPO training with IF-Critic even achieve comparable performance to significantly larger LLMs (32B or 70B) of the same model family. In contrast, the general reward model, Skywork-Reward-V2-Llama-3.1-8B-40M, is not well-suited for instruction-following optimization, yielding only marginal improvements. Secondly, compared to the powerful baseline LLM critic, QwQ-32B, IF-Critic only incurs less than a third of computational overhead, highlighting its practicality. Finally, between the two commonly used optimization methods, DPO and GRPO, GRPO proves more effective for instruction-following tasks, delivering greater performance gains. The reward curves during GRPO training are provided in the Appendix. We also validate that optimization with IF-Critic does not compromise the general performance of LLMs in the Appendix.
Conclusion
@@
In this work, we propose IF-Critic, a LLM critic which is capable of providing fine-grained assessments for instruction-following. IF-Critic adopts a checklist-guided critique generation paradigm that evaluates the following of all constraints in a given checklist within one inference pass, improving its efficiency and reliability. To construct IF-Critic, we collect high-quality critique training data through a multiple-stage critique filtering mechanism and employ a constraint-level preference optimization method to train IF-Critic. Extensive experiments demonstrate that IF-Critic can beat strong LLM-as-a-Judge baselines and bring substantially superior performance gains to LLMs in instruction-following optimization, compared to strong LLM critic baselines.
Appendix A Details of IF-Critic Training

### v2

@@ 3.4 IF-Critic Training
After obtaining the SFT critic, we further enhance its ability via preference optimization. Existing methods usually sample multiple independent responses from the policy model and utilize response-level rewards to select chosen and rejected examples Yuan et al. (2024); Ye et al. (2025b); Yu et al. (2025). However, since the critiques in our task involve multiple segments for different constraints, only the segments where chosen and rejected critiques differ in judgments encapsulate crucial preference information. Other segments may introduce unnecessary interference and dilute important supervision signals, leading to insufficient optimization Chen et al. (2024); Cheng et al. (2025).
To mitigate this, we introduce fine-grained, constraint-level comparisons in preference pairs construction, thus reinforcing the perception of crucial preference information. Specifically, for each evaluation input $p$ in $D_{\text{ref}}$, we first sample $M$ critiques from the SFT critic. Then, we identify the best self-sampled explanation $\hat{e}_{k}$ for each constraint $c_{k}$ whose judgements align with the expert final judgement $j^{*}_{k}$ in a manner analogous to Equation 1 and 2 (if any). Finally, self-sampled critiques with at least one judgment that misaligns with the expert final critique are selected as rejected critiques $C_{l}=\bigcup_{k=1}^{n}(e_{k},j_{k})$. Retaining the segments that align with the expert final critique while replacing other misaligned segments with the best self-sampled explanations $\hat{e}_{k}$ and expert final judgement $j^{*}_{k}$ to construct chosen critiques $C_{w}$:
@@
Rejected critiques are discarded when some misaligned segments lack substitute best self-sampled explanations. Consequently, we construct a preference dataset $D_{\text{dpo}}$ where the preference pairs only differ in the segments with inconsistent judgments, thereby reinforcing crucial preference information. Moreover, self-sampled critiques are closer to the SFT critic decoding space than the expert critique, which is generally beneficial for optimization Guo et al. (2024). Finally, we conduct preference optimization via the following DPO loss:
@@
3.5 Evaluation Results Utilization
@@
4 Experiments
4.1 Instruction-Following Evaluation
@@
Given the scarcity of existing instruction-following meta-evaluation datasets, we sample instructions from four benchmarks: EvalCritic, CFBench Zhang et al. (2024a), TRACE Zhang et al. (2025), and Multi-IF He et al. (2024b) to construct meta-evaluation datasets. EvalCritic is a random split of the dataset we used for instruction-following optimization (§4.2) , whose distribution of instructions differs from IF-Critic training data. For the first three benchmarks, we randomly sample 150 instructions. For Multi-IF, we select the first round from the Chinese version, comprising 454 instructions. Responses for each instruction are generated by 2 of 6 LLMs with varying capabilities. We use our checklist generator to generate constraint checklists for EvalCritic and directly adapt the original checklists for other benchmarks. For the first three benchmarks, two annotators independently judge the following of each constraint, with a third inspector conducting cross-validation. Any discrepancies will be discussed to reach a consensus. For Multi-IF, we use the provided verification code for evaluation. Across these benchmarks, there are 1411 / 1070 / 1634 / 1220 constraints, respectively, of which 406 / 252 / 297 / 320 are labeled as "not followed".
Prompt Model EvalCritic CFBench TRACE Multi-IF Avg.
@@
Constraint-Level Gemini-3-Pro 0.887 0.744 0.934 0.794 0.932 0.704 0.962 0.898 0.857
o4-mini 0.887 0.717 0.927 0.755 0.923 0.599 0.960 0.890 0.832
GPT-4.1 0.883 0.608 0.919 0.689 0.924 0.597 0.932 0.801 0.794
@@
Checklist-Level Gemini-3-Pro 0.902 0.742 0.945 0.808 0.932 0.655 0.960 0.892 0.855
o4-mini 0.911 0.753 0.935 0.761 0.925 0.639 0.964 0.900 0.849
GPT-4.1 0.883 0.560 0.918 0.637 0.923 0.517 0.933 0.799 0.771
@@
We choose state-of-the-art general and evaluation-specific LLMs as our baselines, which are listed in Appendix C. General LLMs can serve as evaluators for instruction-following via elaborately prompting. We explore two prompting strategies: (1) Checklist-Level Prompt: evaluate all constraints in the given checklist within a single inference pass, which is the same as IF-Critic. (2) Constraint-Level Prompt: evaluate each constraint individually. Evaluation-specific LLMs are mainly designed for pairwise comparison. To accommodate this evaluation setting, we convert the above datasets into pairwise versions. Specifically, for two responses generated from the same instruction, we select pairs in which one response follows all constraints while the other does not to construct pairwise samples. The rewards of IF-Critic according to Equation 6 for both responses are converted into pairwise comparisons to calculate the agreement rate. To ensure a fair comparison, we remove samples that result in a tie for IF-Critic. Greedy search decoding is used for reproducibility.
@@
We choose Qwen-2.5-14B-Instruct Yang et al. (2024) as the base model for both the checklist generator and IF-Critic. We also experiment with other base models, and the results are provided in Appendix E. Deepseek-R1 generates $N=5$ expert critiques per response. During IF-Critic training, the ratio of $D_{\text{sft}}$ to $D_{\text{ref}}$ is set to 6:4. The number of self-sampled critiques, $M$, is set to 10, and up to 1 preference pair is constructed for each evaluation input. More implementation details are in Appendix D.
@@
Table 3: The average of Positive F1 and Negative F1 scores under different ablation settings for IF-Critic.
@@
Table 4: Experimental results (%) on Multi-IF, CFBench, and SysBench. The best performance among different critics is bold. Relative Time refers to the relative time to IF-Critic of reward calculation in DPO and average per-step training in GRPO. Skywork-Reward-V2 is a reward model that generates scalar rewards, whereas QwQ-32B and IF-Critic generate natural language critiques.
@@
Table 1 presents the F1-scores for both positive and negative classes in the instruction-following evaluation task of general LLMs, while Table 2 shows the pairwise agreement for evaluation-specific LLMs. Firstly, IF-Critic demonstrates strong performance on various benchmarks, outperforming all baselines on average F1, including o4-mini and Gemini-3-Pro. In contrast, the performance of existing general LLMs remains limited, with relatively low negative F1. This aligns with our motivation for developing IF-Critic to enable more reliable evaluation. Secondly, existing evaluation-specific LLMs still perform poorly in evaluating instruction-following and lag far behind IF-Critic, highlighting the necessity of a critic specifically designed for this task. Finally, reasoning models (e.g, o4-mini, QwQ-32B) typically perform better with Checklist-Level Prompt. This may be attributed that long-chain reasoning enables a more holistic perception of the relationships between constraints with the help of checklists. This observation validates the rationale of our checklist-guided critique generation paradigm.
@@
To investigate the impact of each factor of our framework, we conduct additional ablation studies. For data format, we decompose expert critiques to train an alternative model that generates critiques for individual constraints. Table 3 shows that the performance of IF-Critic declines across all benchmarks, demonstrating that checklist-guided training can facilitate instruction perception and contribute to the final performance.
For critique filtering mechanism, we remove each step individually to explore its impact on evaluation performance. Table 3 shows that all steps contribute to the final performance, but removing rule-augmented verification yields the largest performance drop, underscoring its key role in improving discriminability for length constraints. Notably, directly using all raw training data without critique filtering leads to the most severe performance decline, highlighting its effectiveness.
For preference optimization method, we explored three variants of our method: (1) Constructing preference pairs directly based on the agreement rates between expert critiques, denoted as Vanilla DPO, (2) Replacing misaligned segments of rejected critiques with expert critiques rather than the best self-sampled ones, and (3) Performing SFT on the entire training dataset without preference optimization. Table 3 shows that our constraint-level preference optimization method achieves the best performance, evidencing its effectiveness in enabling better alignment with expert critiques.
@@
To assess the quality of the explanations generated by IF-Critic, we randomly sample 300 constraints from the first three evaluation benchmarks and collect corresponding generated explanations from IF-Critic, QwQ-32B, Deepseek-R1, and o4-mini. For each pair of explanations (one from IF-Critic and the other from a baseline evaluation model, given the same evaluation input), we ask human annotators to judge which explanation is better (i.e., win, lose, or tie) in terms of correctness, informativeness, and helpfulness. The priority of these aspects follows the above order. As shown in Figure 3, IF-Critic surpasses QwQ-32B and Deepseek-R1 in win rates by 9.3% and 7.7%, respectively, and performs on par with o4-mini. Many results are ties because all four evaluators are strong and correct on many constraints. This demonstrates that IF-Critic can generate high-quality explanations for instruction-following evaluation.
4.2 Instruction-Following Optimization
@@
To collect diverse and high-quality complex instructions for model training, we first curate seed instructions from real-world application scenarios. These scenarios are distinct from the training data of IF-Critic and generally exhibit lower instruction complexity. Stratified sampling based on task categories is adopted to enhance diversity. Subsequently, we employ Deepseek-R1 to introduce constraints to these instructions based on a comprehensive constraint taxonomy that encompasses 6 primary constraint categories and 4 constraint composition types. Deepseek-R1 is also used to validate the generated instructions, ensuring they are reasonable, unambiguous, and logically consistent. In total, we obtain 18.5k high-quality complex instructions for training. Further details are in Appendix F.
We choose Qwen2.5-7B-Instruct and Llama-3.1-8B-Instruct as the initial policy models for DPO and GRPO training. Three evaluation models are used to provide reward signals: IF-Critic, the state-of-the-art reward model Skywork-Reward-V2-Llama-3.1-8B-40M, and the powerful open-source general LLM QwQ-32B. For DPO training, we sample 10 responses per instruction, with a temperature of 1.0 and a top-p value of 0.9. The responses with the highest and lowest rewards are used to construct preference pairs. For GRPO training, we perform 32 rollouts per instruction. More training details are in Appendix F.
@@
We evaluate our models on three popular instruction-following benchmarks: Multi-IF He et al. (2024b), which assesses multi-turn and multilingual hard constraints following; CFBench Zhang et al. (2024a), which covers a comprehensive range of constraint types; and Sysbench Qin et al. (2025), which focuses on the following of system prompts. GPT-4o-2024-11-20 is used as the evaluation model for CFBench and SysBench. For Multi-IF, we follow RPO Huang et al. (2025) and report the average performance across the English and Chinese subsets.
@@
The main results are presented in Table 4. Firstly, optimization with IF-Critic brings substantially superior performance gains across various benchmarks compared to all strong baselines, substantiating its advanced capacity of instruction-following evaluation. Notably, LLMs after GRPO training with IF-Critic even achieve comparable performance to significantly larger LLMs (32B or 70B) of the same model family. In contrast, the general reward model, Skywork-Reward-V2, is ill-suited for instruction-following optimization, yielding only marginal improvements. Secondly, compared to the powerful LLM critic, QwQ-32B, IF-Critic only incurs less than a third of computational overhead, highlighting its practicality. Finally, GRPO proves more effective for instruction-following tasks compared to DPO, delivering greater performance gains. The reward curves during GRPO training are provided in the Appendix G. We also validate that optimization with IF-Critic does not compromise the general performance of LLMs in the Appendix H.
5 Conclusion
@@
Our work proposes IF-Critic, an LLM critic for fine-grained instruction-following evaluation. IF-Critic adopts a checklist-guided critique generation paradigm that evaluates all constraints in a checklist within one inference pass, improving efficiency and reliability. We collect high-quality critique training data through a multi-stage filtering mechanism and employ constraint-level preference optimization to train IF-Critic. Extensive experiments show that IF-Critic has stronger evaluation ability and can bring superior performance gains to LLMs in instruction-following optimization, compared to powerful LLM critic baselines.
6 Limitations
The limitations of our work are summarized as follows:
Scope of Rule-Augmented Verification.
In the critique training data construction stage of IF-Critic, we only introduce rule-augmented verification for length-related constraints, while other code-verifiable constraints (e.g., keywords or structural format) are not comprehensively covered. We prioritize length-related constraints because they are prevalent in real-world complex instructions, and the previous version of IF-Critic performs well on other verifiable constraints but struggles with length constraints, as observed in our primary experiments. Although this approach has yielded substantial performance gains, extending this approach to a broader spectrum of constraint types remains an important direction for future work.
Potential Evaluation Bias.
Similar to other model-based evaluation methods, IF-Critic may also suffer from potential evaluation biases such as self-enhancement or verbosity bias Zheng et al. (2023). These biases could harm the evaluation correctness of certain models. Although our multi-stage critique filtering mechanism has effectively mitigated noise and bias in training critique collection, incorporating inference-time strategies (e.g., multi-agent debate) may further attenuate bias and enhance evaluation performance. We reserve further investigation of evaluation bias mitigation as important future work.
Ethical Considerations
In this work, we recruit some human annotators to validate the quality of training data and the performance of IF-Critic. The annotator pool primarily consists of college students. Throughout the data annotation process, we adhere to several key principles. Firstly, all annotators are fully informed about the purpose of our study and the involved tasks. Secondly, all annotators received fair compensation for their contributions based on the market price. Finally, any harmful instructions are filtered before annotation to avoid ethical issues.
Appendix A Model List for Response Generation
We use 15 representative LLMs for response generation, including GPT-4o OpenAI (2023), GPT-4o-mini OpenAI (2023), seven versions of GLM-4 GLM et al. (2024), three versions of Qwen2.5 Yang et al. (2024), Deepseek-R1 Guo et al. (2025a), Doubao1.5 Doubao Team (2025), and Claude3.7-sonnet Anthropic (2023).
Appendix B List for Prompt Templates
This section lists all the prompt templates applied throughout this work, including the prompt for scoring the quality of user instruction in Table 8, the prompt for constraint checklist generation in Table 9, the prompt for checklist-guided critique generation in Table 10, the prompt for cross-model verification in Table 11, and the prompt for rule-augmented verification in Table 12 and 13.
Appendix C List for Baseline Evaluation Models
For general LLMs, we adopt 3 advanced proprietary LLMs, GPT-4.1 OpenAI (2023), o4-mini Jaech et al. (2024), and Gemini-3-Pro Team et al. (2023), as well as 5 strong open-source LLMs, GLM-4-32B-0414 GLM et al. (2024), LLama-3.3-70B-Instruct Grattafiori et al. (2024), Qwen2.5-72B-Instruct Yang et al. (2024), QwQ-32B Team (2025), and Deepseek-R1 Guo et al. (2025a) as our general baselines. For evaluation-specific LLMs, we select 4 powerful discriminative or generative reward models, Skywork-Reward-V2-Llama-3.1-8B-40M Liu et al. (2025a), Prometheus-BGB-8x7B Kim et al. (2024b), RM-R1-Deepseek-Distilled-Qwen-32B Chen et al. (2025), and RRM-32B Guo et al. (2025b) as our evaluation-specific baselines.
Appendix D Implementation Details of IF-Critic